# Sum of Two Squares in Biquadratic Fields

### v3

@@ Introduction
Some critical values about $F$, including integers in $F$ and discriminant of $F$, is already given in [1].
In this article, using local-global principle of isotropy of quadratic forms (see 66:1 of [2]), we will give an explicit algorithm to determine whether a number $s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}$ ($s_{0},s_{1},s_{2},s_{3}\in\mathbb{Z}$ and coprime) is a sum of two squares in $\mathbb{F}$. First we prepare some lemmas for convenience:
@@ 1 Prime Numbers
First we state some facts. For any odd prime number $p$, either $p\nmid abc$, or $p$ exactly divides two of $\{a,b,c\}$. In the former case $(\frac{a}{p})(\frac{b}{p})=(\frac{c}{p})$ ($**\frac{*}{*}$ denotes Legendre symbol), and in the latter case, say $p|a$ and $p|b$, $(\frac{a/p}{p})(\frac{b/p}{p})=(\frac{c}{p})$.
We already know that a prime number $q$ is a sum of two squares in $\mathbb{Q}$ if and only if it $q\not\equiv 3\pmod{4}$. Let $q$ be a prime number with $q\equiv 3(mod4)$. Next we find all biquadratic fields where $q$ becomes a sum of two squares.
@@
(2)Of course $(\frac{-1,q}{\mathcal{P}})=1$ at infinite places. We only need to discuss $q$-adic and dyadic cases. By lemma 0.2, $(\frac{-1,q}{\mathcal{P}})=1$ if and only if $2O_{F}$ does not completely split. In this case, if $(\frac{-1,q}{\mathcal{Q}})=-1$ for $q$-adic place $\mathcal{Q}$, then $F_{\mathcal{Q}}$ must not contain $\mathbb{Q}_{q}(\sqrt{-1})$, $\mathbb{Q}_{q}(\sqrt{q})$ or $\mathbb{Q}_{q}(\sqrt{-q})$ (in the last case $(\frac{-1,-1}{\mathcal{Q}})=-1$ , making $(\frac{-1,q}{\mathcal{Q}})=(\frac{-1,-1}{\mathcal{Q}})(\frac{-1,-q}{\mathcal{Q}})=1$). So $F_{\mathcal{Q}}\simeq\mathbb{Q}_{q}$, i.e, $(\frac{a}{q})=(\frac{b}{q})=1$. ∎
@@ 2 General cases
Let $S=s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}\in F$ in introduction. By 66:1 and 63:12 of [2], to decide whether $S$ is a sum of squares in $F$, we only need to calculate $(\frac{-1,M}{\mathcal{P}})$s for all infinite, dyadic or $p$-adic places for $p\equiv 3\pmod{4}$ and $p|N_{F/\mathbb{Q}}(M)$. The infinite case is obvious: $(\frac{-1,S}{\mathcal{\infty}})=-1$ if and only if $F$ is real and $S$ is not totally positive.
Let $\mathcal{P}\cap\mathbb{Z}=p\equiv 3(mod4)$, $p|N_{F/\mathbb{Q}}(S)$.
@@ 2.1 Algorithm about unramified $p$
Therefore given an algorithm to calculate powers of a number in an ideal of a quadratic field, that in a biquadratic field can be obtained.
@@
If $p\equiv 3(mod4)$ is a prime dividing $N_{F/\mathbb{Q}}(S)$ with $(\frac{a}{p})=(\frac{b}{p})=(\frac{c}{p})=1$, then $(\frac{-1,S}{\mathcal{P}})=1$ for all for $p$-adic places $\mathcal{P}$, if and only if
@@ 2.2 Algorithm about ramified $p$
If $p\equiv 3(mod4)$ is a prime dividing $N_{F/\mathbb{Q}}(S)$, and $abc$ (say, $p|a$ and $p|b$), and $(\frac{a/p}{p})=(\frac{b/p}{p})$ (equivalent that $(\frac{c}{p})=1$), then $(\frac{-1,S}{\mathcal{P}})=1$ for all for $p$-adic places $\mathcal{P}$, if and only if both $v_{\mathfrak{p}_{1}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{2}}(S\sigma_{2}(S))-2v_{\mathfrak{p}_{3i}}(\sigma_{1}(S)\sigma_{2}(S))$ are divided by 4 for $i\in\{1,2\}$.
@@ 2.3 Dyadic cases
(2)If $b\equiv 5\pmod{8}$, i.e, $2O_{F}$ splits and inerts, then both $F_{\mathcal{P}}$s are isomorphic to $\mathbb{Q}_{2}(\sqrt{5})$. If $S$ is a sum of two squares in $F$, then let $S^{\prime\prime}=\frac{S}{2^{v_{2}(S)}}$($0\leq v_{2}(S)\leq 2$) and

### v4

@@ Introduction
Some critical values about $F$, including integers in $F$ and discriminant of $F$, are already given in [1].
In this article, using local-global principle of isotropy of quadratic forms (see 66:1 of [2]), we will give an explicit algorithm to determine whether a number $s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}$ ($s_{0},s_{1},s_{2},s_{3}\in\mathbb{Z}$) is a sum of two squares in ${F}$. First we prepare some lemmas for convenience (see [4]):
@@ 1 Prime Numbers
First we state some facts. For any odd prime number $p$, either $p\nmid abc$, or $p$ exactly divides two of $\{a,b,c\}$. In the former case $(\frac{a}{p})(\frac{b}{p})=(\frac{c}{p})$ ($\frac{*}{*}$ denotes Legendre symbol), and in the latter case, say $p|a$ and $p|b$, $(\frac{a/p}{p})(\frac{b/p}{p})=(\frac{c}{p})$.
It is well-known that a prime number $q$ is a sum of two squares in $\mathbb{Q}$ if and only if it $q\not\equiv 3\pmod{4}$. Let $q$ be a prime number with $q\equiv 3\pmod{4}$. Next we find all biquadratic fields where $q$ becomes a sum of two squares.
@@
(2)Of course $(\frac{-1,q}{\mathcal{P}})=1$ at infinite places. We only need to discuss $q$-adic and dyadic cases. By lemma 0.2, $(\frac{-1,q}{\mathcal{P}})=1$($\mathcal{P}$ dyadic) if and only if $2O_{F}$ does not completely split. In this case, if $(\frac{-1,q}{\mathcal{Q}})=-1$ for $q$-adic place $\mathcal{Q}$, then $F_{\mathcal{Q}}$ must not contain $\mathbb{Q}_{q}(\sqrt{-1})$, $\mathbb{Q}_{q}(\sqrt{q})$ or $\mathbb{Q}_{q}(\sqrt{-q})$ (in the last case $(\frac{-1,-1}{\mathcal{Q}})=1$ , making $(\frac{-1,q}{\mathcal{Q}})=(\frac{-1,-1}{\mathcal{Q}})(\frac{-1,-q}{\mathcal{Q}})=1$). So $F_{\mathcal{Q}}\simeq\mathbb{Q}_{q}$, i.e, $(\frac{a}{q})=(\frac{b}{q})=1$. ∎
@@ 2 General cases
Let $S=s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}\in F$ in introduction. By 66:1 and 63:12 of [2], to decide whether $S$ is a sum of squares in $F$, we only need to calculate $(\frac{-1,S}{\mathcal{P}})$s for all infinite, dyadic or $p$-adic places for $p\equiv 3\pmod{4}$ and $p|N_{F/\mathbb{Q}}(S)$. The infinite case is obvious: $(\frac{-1,S}{\mathcal{\infty}})=-1$ if and only if $F$ is real and $S$ is not totally positive.
Let $\mathcal{P}\cap\mathbb{Z}=p\equiv 3\mod 4$, $p|N_{F/\mathbb{Q}}(S)$.
@@ 2.1 Algorithm about unramified $p$
Therefore given an algorithm to calculate powers of a number in a prime ideal of a quadratic field, that in a biquadratic field can be obtained.
@@
If $p\equiv 3\pmod{4}$ is a prime dividing $N_{F/\mathbb{Q}}(S)$ with $(\frac{a}{p})=(\frac{b}{p})=(\frac{c}{p})=1$, then $(\frac{-1,S}{\mathcal{P}})=1$ for all four $p$-adic places $\mathcal{P}$, if and only if
@@ 2.2 Algorithm about ramified $p$
If $p\equiv 3\pmod{4}$ is a prime dividing $N_{F/\mathbb{Q}}(S)$, and $abc$ (say, $p|a$ and $p|b$), and $(\frac{a/p}{p})=(\frac{b/p}{p})$ (equivalent that $(\frac{c}{p})=1$), then $(\frac{-1,S}{\mathcal{P}})=1$ for both $p$-adic places $\mathcal{P}$, if and only if both $v_{\mathfrak{p}_{1}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{2}}(S\sigma_{2}(S))-2v_{\mathfrak{p}_{3i}}(\sigma_{1}(S)\sigma_{2}(S))$ are divided by 4 for $i\in\{1,2\}$.
@@ 2.3 Dyadic cases
(2)If $b\equiv 5\pmod{8}$, i.e, $2O_{F}$ splits and inerts, then both $F_{\mathcal{P}}$s are isomorphic to $\mathbb{Q}_{2}(\sqrt{5})$. Let $S^{\prime\prime}=\frac{S}{2^{v_{2}(S)}}$($0\leq v_{2}(S)\leq 2$) and
@@
Acknowledgements
The author is sincerely grateful for directions by Prof Hourong Qin, and support by National Natural Science Foundation of China (11971224).